# A galactic outflow traced by its extended Mg II emission out to a $\sim 30$ kpc radius in the Hubble Ultra Deep Field with MUSE

### v1

@@
(Received XXX; Accepted XXX)
@@ 1 Introduction
Observing the large-scale gaseous environment of galaxies is a major challenge, because of the large projected solid angles covered by galactic haloes and the low densities of the material within (Anglés-Alcázar et al., 2017; Tumlinson et al., 2017; Augustin et al., 2019; Corlies et al., 2020). Most constraints come from absorption line studies in privileged lines of sight probed by sufficiently bright background objects, usually quasars (see, e.g., Bergeron & Stasińska, 1986; Werk et al., 2013). These have brought up the notion of a generic circumgalactic medium (CGM), where several processes compete to feed the galaxies or drive gas out of them (see Faucher-Giguère & Oh, 2023, for a review on the topic). The spatial distribution of the CGM has been studied mainly through the statistical combination of many independent absorption sightlines. One of the key results of the last years is the clear bimodality of strong low-ionization metal absorption, in particular, the Mg II $\lambda\lambda$2796, 2803 doublet, with respect to the symmetry axes of the central galaxies, indicating a significantly non-isotropic distribution of the circumgalactic gas (e.g., Bordoloi et al., 2011; Bouché et al., 2012; Schroetter et al., 2019), in agreement with results from cosmological simulations (e.g., Péroux et al., 2020). For $L^{\star}$ galaxies the data are consistent with two main components of the absorbing CGM: A cone-like outflow in the polar direction, and an extended disc-like structure close to the equatorial plane (Bouché et al., 2013; Ho et al., 2017; Zabl et al., 2019), plausibly related to ongoing gas accretion from the intergalactic medium (although findings from Pointon et al., 2019, suggest that cold streams of accreting filaments are not necessarily aligned with the galaxy plane).
Structural information on individual objects, rather than ensemble averages, is however very hard to obtain. With the absorption line technique, this is restricted to rare cases where multiple absorption lines of sight are available for a single galaxy. Even this ‘tomographic’ approach can only provide a very sparse sampling of any particular CGM region. Such measurements have been used to constrain the amount of gas clumping, turbulence, and metallicities on scales of up to kpc (e.g., Smette et al., 1995; Lopez et al., 1999; Chen et al., 2014; Zahedy et al., 2016; Lopez et al., 2018; Rubin et al., 2018; Peroux et al., 2018; Afruni et al., 2023; Dutta et al., 2024), and also to test the wind model for Mg II absorbers (Zabl et al., 2020).
@@
The first detection of extended Mg II emission around a galaxy was reported by Rubin et al. (2011), who found excess emission over the continuum around a galaxy at $z=0.69$ out to $\sim 10$ kpc on both sides. The second and quite similar object, a $z=0.94$ galaxy with a one-sided Mg II extension of again $\sim 10$ kpc (Martin et al., 2013), was however identified as the only such case out of a sample of 145 spectroscopically observed galaxies with coverage of Mg II. In addition to these two individual detections, Erb et al. (2012) presented statistical evidence that Mg II emission in a stack of longslit spectra of star-forming galaxies at $z=1.4$–2 is slightly more extended than the continuum, although the detected excess is very modest. While these studies were hampered by the inevitable limitations of longslit spectroscopy, inhibiting a complete picture of the distribution of Mg II emission around the observed galaxies, they suggest that circumgalactic Mg II emission is either a rare phenomenon or very hard to detect (or both). The non-detection of any extended emission around five galaxies in deep narrow band imaging by Rickards Vaught et al. (2019) is consistent with this notion.
@@
In a more recent paper, Burchett et al. (2021) present spatially-resolved observations from the Keck Cosmic Web Imager (KCWI) of the same galaxy studied by Rubin et al. (2011). Using integral field spectroscopic data, the authors are able to observe in 3D the circumgalactic Mg II emission around the central galaxy, and put some constraints on the physical properties of the CGM. Comparing their data with the radiative transfer models introduced by Prochaska et al. (2011), they find it to be consistent with a radially decelerating isotropic wind. This represents the first spatially resolved modeling of a Mg II halo using integral field spectroscopic data.
@@
In a qualitatively similar study, using data from the Multi-Unit Spectroscopic Explorer (MUSE; Bacon et al., 2010), Zabl et al. (2021) report the first detection of a Mg II halo in emission, up to distances of 25 kpc from the central galaxy, that is simultaneously probed in absorption by the sightline of a bright background QSO at an impact parameter of $\sim 40$ kpc. The authors find good agreement between the kinematics of the observed emission and absorption, with those produced by a simple toy model of a biconical outflow with a constant velocity, in contrast with the radially decelerating isotropic wind model from Burchett et al. (2021).
@@
Leclercq et al. (2022) report the discovery of extended Mg II emission, emerging from the intragroup medium of a group of five star-forming galaxies at $z=1.31$, with a maximal projected extent of $\sim 70$ kpc. They also find evidence suggesting this gas has been ejected through stellar outflows from these galaxies towards the intragroup medium. Leclercq et al. (2024) also find evidence of extended Mg II (and [O II]) emission in a sample of confirmed Lyman continuum leakers and non-leakers at $z\approx 0.35$, aiming to connect the neutral and ionized gas distributions to the Lyman continuum escape fraction.
@@
The work by Burchett et al. (2021), Zabl et al. (2021), Leclercq et al. (2022), and Leclercq et al. (2024) demonstrates how much we can learn about the galactic wind properties through a fully spatially resolved analysis. However, there is still no clear consensus regarding the general properties of galactic winds, such as their velocity, acceleration, precise geometry, outflow mass rate, or loading factor. On the other hand, this kind of analysis requires very deep, high-S/N data, and as a consequence, there is only a limited number of examples, insufficient to provide a representative view of the CGM in emission. It is clearly necessary to push further in the direction of spatially resolved analyses that preserve the overall structure of the CGM, towards a more robust understanding of the mechanisms that shape it.
@@
In this paper, we present the discovery and analysis of the Mg II emission halo around a galaxy at $z=0.737$, obtained in observations with the panoramic imaging spectrograph MUSE. We use the high sensitivity and spatial resolution of MUSE to detect Mg II emission up to distances of more than 20 effective radii from the central galaxy, and then we use a simple outflow modeling scheme to model the inner regions of the halo, allowing us to interpret the observations and put constraints on the properties of galactic outflows. The paper is structured as follows: We summarize our data in Sec. 2, and give an overview of the properties of our studied galaxy in Sec. 3. In Sec. 4 we emphasize the spatially resolved characteristics of the Mg II emission and absorption in this galaxy. In Sec. 5, we introduce the model used to interpret our observations, and in Sec. 6, we describe the methodology to fit our data. We describe our modeling results in Sec. 7, and in Sec. 8, we discuss them, together with the limitations of our approach, as well as putting our results in the context of previous findings. Finally, in Sec. 9, we summarize our results and present some conclusions and perspectives. We use a flat concordance cosmological model with $H_{0}=68$ km s-1 Mpc-1 and $\Omega_{\mathrm{m}}=0.3$. All quoted transverse distances are in physical units, not comoving.
@@ 2 Observational data
This study is based on observations with the MUSE instrument on the ESO-VLT. MUSE is a panoramic integral-field spectrograph with a field of fiew of $1^{\prime}\times 1^{\prime}$ (in Wide Field Mode, used for these observations) at a spatial sampling of $0\aas@@fstack{\prime\prime}2\times 0\aas@@fstack{\prime\prime}2$. The spectral range is 4750 Å–9350 Å, with a spectral sampling of 1.25 Å and a wavelength-dependent velocity resolution between $\sim$190 and $\sim$80 km s-1 (FWHM) at the blue and red ends of the spectrum, respectively.
@@
Between 2014 and 2016 we obtained a $3^{\prime}\times 3^{\prime}$ MUSE mosaic in the Hubble Ultra Deep Field (HUDF). The effective integration time varies slightly with the position due to the dithering and coadding process; for the region of interest in this paper it is 9.94 hours. The observations and subsequent data processing of the MUSE Mosaic in the HUDF are described in detail in Bacon et al. (2017) and Bacon et al. (2023). The data are seeing-limited, with a resolution (FWHM of the best-fitting Moffat (1969) function) of ${\prime\prime}$63 at 7000 Å (${\prime\prime}$71 at 4850 Å, the wavelength of the Mg II line in the main galaxy of this study). A first catalogue with identifications and redshift measurements of 1439 objects in the MUSE-HUDF mosaic was published in Inami et al. (2017). The data has then been reprocessed with improved flatfielding and sky subtraction, followed by a second pass through the entire dataset using several new or updated tools; the outcome of this reprocessing has been published in Bacon et al. (2023) as part of the second data release of the MUSE-HUDF [^1] Publicly available for download in the AMUSED website https://amused.univ-lyon1.fr/.. The natural coordinate system of the MUSE datacube is tilted by $-44$° with respect to the North-South direction.
@@
In addition to the MUSE observations, we employed the extensive multiwavelength data available for the HUDF via public archives, in particular the deep optical and near-infrared images obtained with the Hubble Space Telescope (Beckwith et al., 2006; Illingworth et al., 2013; Oesch et al., 2018), mid-infrared photometry by the Spitzer Space Telescope (Labbé et al., 2015), and X-ray observations of the Chandra satellite (Luo et al., 2017).
@@
Figure 1: HST/ACS image of the galaxy MUSE-UDF #884 (in the field centre) and its environment in the F775W filter. The central parts of the brightest galaxies are shown at higher cut levels in yellow. The inset in the lower left provides an enlarged view of the image centre indicated by the white square. Spectroscopic redshifts measured with MUSE are labelled in black, always referring to the object left of the label. The gray labels give photometric redshift ranges taken from Rafelski et al. (2015), estimated with the BPZ code. Galaxies mentioned in the text are marked as ‘NB’ (physical neighbour), ’FG’ (foreground) and ’BG (background). The image orientation follows that of the MUSE datacube and is given by the compass marker in the lower right.
@@ 3.1 Discovery and environment
As part of the above-mentioned reprocessing of the MUSE-HUDF Mosaic data, we searched again for emission line peaks in the datacube, using the LSDCat software (Line Source Detection and Cataloguing, Herenz & Wisotzki, 2017). This blind search revealed two significant peaks, above our signal-to-noise (S/N) detection threshold of 5.0, that turned out to be a partial ring of Mg II $\lambda$2796 emission (with its corresponding Mg II $\lambda$2803 counterpart) surrounding a previously known galaxy at $z=0.737$. At the redshift of this galaxy, 1″ corresponds to $7.5$ kpc.
@@
The central galaxy itself has cross-identifications222e.g., ID #23061 in UVUDF (Rafelski et al., 2015), ID #13337 in CANDELS (Guo et al., 2013), ID #25592 in 3D-HST (Skelton et al., 2014) in many surveys of the HUDF/CDFS region. It is listed as ID #884 in the catalogue of the MUSE-HUDF mosaic by Inami et al. (2017), and for brevity we will refer to it by only this number in the following. It is a luminous blue galaxy with $m_{\mathrm{AB}}=21.7$ in the ACS/F775W band, implying $L\approx L^{\star}$ at $z=0.737$. Figure 1 shows a deep HST image of the galaxy and its immediate surroundings within ±$\pm 11\arcsec$ ($\pm$80 kpc). Several of the galaxies close to the line of sight have redshifts measured with MUSE and are included in the Inami et al. (2017) sample; these are labelled in the figure. We also provide a few photometric redshifts (taken from Rafelski et al., 2015) where this may be of interest. Galaxies unlabelled in Fig. 1 have no reliable redshift information or are most likely unrelated to #884. The physically closest galaxy – marked as ‘NB1’ – is MUSE-HUDF #6516 ($m_{\mathrm{AB}}=23.8$) at 53 kpc projected distance and with only $-85$ km s-1 redshift difference. Other than this, there are no surrounding objects that are plausibly located nearby in real space; the galaxy at $z=0.734$ near the bottom of the field has a redshift separation of $-1000$ km s-1 and is thus probably already in the foreground. On transverse scales of up to $\sim 1$ Mpc, the full MUSE-HUDF Mosaic reveals 23 (11) galaxies within $\pm$ 1000 (300) km s-1, most of them faint emission line objects scattered all over the field. There is no evidence for #884 to be located in a cluster or rich group environment.
@@
Other projected neighbors to #884 are marked in Fig. 1 as ‘BG’ (background) and ‘FG’ (foreground). Since these galaxies are faint, we used the TDOSE software package (Three-Dimensional Optimal Spectral Extraction, Schmidt et al., 2019) to perform an optimal spectra extraction. Each of the galaxies was modelled in the HST data as a superposition of multiple components, which after blurring to the MUSE PSF were then used to define the spatial profiles of the galaxies used in the extraction. We briefly comment on three of these projected neighbors:
@@
  • BG1 (= MUSE-HUDF #6943 in Inami et al. 2017) is a spectroscopically confirmed system at $z=1.221$, at 4″ distance to #884, and with $m_{\mathrm{AB,F775W}}=24.5$ it is sufficiently bright to produce a well-detected continuum in the MUSE spectrum. The continuum-subtracted datacube shows two significantly negative peaks – indicative of absorption in the subtracted continuum – at the location of BG1 and at the wavelengths of the Mg II emission around #884. We analyze and discuss this interesting sightline in Sec. 8.3.
@@
  • The galaxy marked as FG1 ($m_{\mathrm{AB,F775W}}=25.3$) has a projected distance of only 2″ to #884, but no redshift in Inami et al. (2017). The photometric redshift estimate in Rafelski et al. (2015) puts it in the foreground, $0.530.38<z_{\mathrm{BPZ}}<$ 0.53. The extracted MUSE spectrum does not show any emission features, apart from [O II] spilling over from #884, and it is too noisy to allow for a confident identification of individual absorption lines. It however agrees well in shape with a synthetic galaxy spectrum from the Jimenez et al. (2004) spectral library, for an age of $\sim 2$ Gyr at a redshift of $z=0.621\pm 0.002$. This redshift is outside the range estimated for the photometric redshift, however, assessing the reliability of the Rafelski et al. (2015) catalog is outside the scope of this paper. While our redshift measurement is not as secure as those based on multiple individual spectral lines, we consider it sufficiently robust to conclude that this galaxy is indeed located in the foreground to #884 and thus cannot provide another absorption line of sight.
    While this redshift is outside the range estimated for the photometric redshift, assesing the reliability of the Rafelski et al. (2015) is beyond the scope of this paper. However, in both redshift estimations put

### v3

@@
(Received 29 April 2024; Accepted 28 August 2024)
@@ 1 Introduction
Observing the large-scale gaseous environment of galaxies is a major challenge because of the large projected solid angles covered by galactic haloes and the low densities of the material within (Anglés-Alcázar et al., 2017; Tumlinson et al., 2017; Augustin et al., 2019; Corlies et al., 2020). Most constraints come from absorption line studies in privileged lines of sight probed by sufficiently bright background objects, usually quasars (see, e.g., Bergeron & Stasińska, 1986; Werk et al., 2013). These have brought up the notion of a generic circumgalactic medium (CGM) where several processes compete to feed galaxies or drive gas out of them (see Faucher-Giguère & Oh, 2023, for a review on the topic). The spatial distribution of the CGM has been studied mainly through the statistical combination of many independent absorption sight lines. One of the key results of the past years is the clear bimodality of strong low-ionization metal absorption, in particular, the Mg II $\lambda\lambda$2796, 2803 doublet, with respect to the symmetry axes of the central galaxies, as it indicates a significantly non-isotropic distribution of the circumgalactic gas (e.g., Bordoloi et al., 2011; Bouché et al., 2012; Schroetter et al., 2019), in agreement with results from cosmological simulations (e.g., Péroux et al., 2020). For $L^{\star}$ galaxies, the data are consistent with two main components of the absorbing CGM: a cone-like outflow in the polar direction and an extended disc-like structure close to the equatorial plane (Bouché et al., 2013; Ho et al., 2017; Zabl et al., 2019) plausibly related to ongoing gas accretion from the intergalactic medium (although findings from Pointon et al., 2019, suggest that cold streams of accreting filaments are not necessarily aligned with the galaxy plane).
However, structural information on individual objects, rather than ensemble averages, is very hard to obtain. With the absorption line technique, this is restricted to rare cases where multiple absorption lines of sight are available for a single galaxy. Even this ”tomographic” approach can only provide a very sparse sampling of any particular CGM region. Such measurements have been used to constrain the amount of gas clumping, turbulence, and metallicities on scales of up to a few kiloparsecs (e.g., Smette et al., 1995; Lopez et al., 1999; Chen et al., 2014; Zahedy et al., 2016; Lopez et al., 2018; Rubin et al., 2018; Peroux et al., 2018; Afruni et al., 2023; Dutta et al., 2024) and to test the wind model for Mg II absorbers (Zabl et al., 2020).
@@
The first detection of extended Mg II emission around a galaxy was reported by Rubin et al. (2011), who found excess emission over the continuum around a galaxy at $z=0.69$ out to $\sim 10$ kpc on both sides. The second and quite similar object, a $z=0.94$ galaxy with a one-sided Mg II extension of again $\sim 10$ kpc (Martin et al., 2013), was identified as the only such case out of a sample of 145 spectroscopically observed galaxies with coverage of Mg II. In addition to these two individual detections, Erb et al. (2012) presented statistical evidence that Mg II emission in a stack of long-slit spectra of star-forming galaxies at $z=1.4$–2 is slightly more extended than the continuum, although the detected excess is very modest. While these studies were hampered by the inevitable limitations of long-slit spectroscopy, inhibiting a complete picture of the distribution of Mg II emission around the observed galaxies, they suggest that ”circumgalactic” Mg II emission is either a rare phenomenon or very hard to detect (or both). The non-detection of any extended emission around five galaxies in deep narrow-band imaging by Rickards Vaught et al. (2019) is consistent with this notion.
@@
In a more recent paper, Burchett et al. (2021) have presented spatially resolved observations from the Keck Cosmic Web Imager (KCWI) of the same galaxy studied by Rubin et al. (2011). Using integral field spectroscopic data, the authors were able to observe in 3D the circumgalactic Mg II emission around the central galaxy, and they put some constraints on the physical properties of the CGM. Comparing their data with the radiative transfer models introduced by Prochaska et al. (2011), they find it to be consistent with a radially decelerating isotropic wind. This represents the first spatially resolved modeling of a Mg II halo using integral field spectroscopic data.
@@
In a qualitatively similar study using data from the Multi-Unit Spectroscopic Explorer (MUSE; Bacon et al., 2010), Zabl et al. (2021) reported the first detection of a Mg II halo in emission, up to distances of 25 kpc from the central galaxy, that is simultaneously probed in absorption by the sight line of a bright background quasi-stellar object (QSO) at an impact parameter of $\sim 40$ kpc. The authors find good agreement between the kinematics of the observed emission and absorption, with those produced by a simple toy model of a biconical outflow having a constant velocity, in contrast to the radially decelerating isotropic wind model from Burchett et al. (2021).
@@
Leclercq et al. (2022) report the discovery of extended Mg II emission emerging from the intragroup medium of a group of five star-forming galaxies at $z=1.31$, with a maximal projected extent of $\sim 70$ kpc. They also find evidence suggesting this gas has been ejected through stellar outflows from these galaxies toward the intragroup medium. Moreover, Leclercq et al. (2024) have found evidence of extended Mg II (and [O II]) emission in a sample of confirmed Lyman continuum leakers and non-leakers at $z\approx 0.35$, aiming to connect the neutral and ionized gas distributions to the Lyman continuum escape fraction.
@@
The work by Burchett et al. (2021), Zabl et al. (2021), Leclercq et al. (2022), and Leclercq et al. (2024) demonstrates how much we can learn about the galactic wind properties through a fully spatially resolved analysis. However, there is still no clear consensus regarding the general properties of galactic winds, such as their velocity, acceleration, precise geometry, outflow mass rate, or loading factor. On the other hand, this kind of analysis requires very deep, high S/N data, and as a consequence, there is only a limited number of examples, which is insufficient to provide a representative view of the CGM in emission. It is clearly necessary to push further in the direction of spatially resolved analyses that preserve the overall structure of the CGM and toward a more robust understanding of the mechanisms that shape it.
@@
In this paper, we present the discovery and analysis of the Mg II emission halo around a galaxy at $z=0.737$ obtained in observations with the panoramic imaging spectrograph MUSE. We used the high sensitivity and spatial resolution of MUSE to detect Mg II emission up to distances of more than 20 effective radii from the central galaxy, and then we used a simple outflow modeling scheme to model the inner regions of the halo, allowing us to interpret the observations and put constraints on the properties of galactic outflows. The paper is structured as follows: We summarize our data in Sect. 2, and we give an overview of the properties of our studied galaxy in Sect. 3. In Sect. 4, we emphasize the spatially resolved characteristics of the Mg II emission and absorption in this galaxy. In Sect. 5, we introduce the model used to interpret our observations, and in Sect. 6, we describe the methodology to fit our data. We describe our modeling results in Sect. 7, and in Sect. 8, we discuss them together with the limitations of our approach and put our results in the context of previous findings. Finally, in Sect. 9, we summarize our results and present some conclusions and perspectives. We used a flat concordance cosmological model with $H_{0}=68$ km s-1 Mpc-1 and $\Omega_{\mathrm{m}}=0.3$. All quoted transverse distances are in physical units, not co-moving.
@@ 2 Observational data
This study is based on observations with the MUSE instrument on the ESO-VLT. MUSE is a panoramic integral-field spectrograph with a field of view (FoV) of $1^{\prime}\times 1^{\prime}$ (in Wide Field Mode, used for these observations) at a spatial sampling of $0\aas@@fstack{\prime\prime}2\times 0\aas@@fstack{\prime\prime}2$. The spectral range is 4750 Å–9350 Å, with a spectral sampling of 1.25 Å and a wavelength-dependent velocity resolution between $\sim$190 and $\sim$80 km s-1 (full width at half maximum; FWHM) at the blue and red ends of the spectrum, respectively.
@@
Between 2014 and 2016 we obtained a $3^{\prime}\times 3^{\prime}$ MUSE mosaic in the Hubble Ultra Deep Field (HUDF). The effective integration time varies slightly with the position due to the dithering and coadding process; for the region of interest in this paper it is 9.94 hours. The observations and subsequent data processing of the MUSE Mosaic in the HUDF are described in detail in Bacon et al. (2017) and Bacon et al. (2023). The data are seeing-limited, with a resolution (FWHM of the best-fitting Moffat (1969) function) of ${\prime\prime}$63 at 7000 Å (${\prime\prime}$71 at 4850 Å, the wavelength of the Mg II line in the main galaxy of this study). A first catalog with identifications and redshift measurements of 1439 objects in the MUSE-HUDF mosaic was published in Inami et al. (2017). The data has then been reprocessed with improved flat fielding and sky subtraction, followed by a second pass through the entire dataset using several new or updated tools; the outcome of this reprocessing has been published in Bacon et al. (2023) as part of the second data release of the MUSE-HUDF [^1] Publicly available for download in the AMUSED website https://amused.univ-lyon1.fr/.. The natural coordinate system of the MUSE data cube is tilted by $-44$° with respect to the north-south direction.
@@
In addition to the MUSE observations, we employed the extensive multiwavelength data available for the HUDF via public archives, in particular the deep optical and near-infrared images obtained with the Hubble Space Telescope (HST; Beckwith et al., 2006; Illingworth et al., 2013; Oesch et al., 2018), mid-infrared photometry by the Spitzer Space Telescope (Labbé et al., 2015), and X-ray observations of the Chandra satellite (Luo et al., 2017).
@@
Figure 1: Image from HST/ACS of the galaxy MUSE-UDF #884 (in the field center) and its environment in the F775W filter. The central parts of the brightest galaxies are shown at higher cut levels in yellow. The inset in the lower left provides an enlarged view of the image center, indicated by the white square. Spectroscopic redshifts measured with MUSE are labeled in black, always referring to the object left of the label. The gray labels give photometric redshift ranges taken from Rafelski et al. (2015), estimated with the BPZ code. Galaxies mentioned in the text are marked as ‘NB’ (physical neighbor), ‘FG’ (foreground) and ‘BG’ (background). The image orientation follows that of the MUSE data cube and is given by the compass marker on the lower right.
@@ 3.1 Discovery and environment
As part of the above-mentioned reprocessing of the MUSE-HUDF Mosaic data, we searched again for emission line peaks in the data cube, using the LSDCat software (Line Source Detection and Cataloguing, Herenz & Wisotzki, 2017). This blind search revealed two significant peaks, above our signal-to-noise (S/N) detection threshold of 5.0, which turned out to be a partial ring of Mg II $\lambda$2796 emission (with its corresponding Mg II $\lambda$2803 counterpart) surrounding a previously known galaxy at $z=0.737$. At the redshift of this galaxy, 1″ corresponds to $7.5$ kpc.
@@
The central galaxy itself has cross-identifications222e.g., ID #23061 in UVUDF (Rafelski et al., 2015), ID #13337 in CANDELS (Guo et al., 2013), ID #25592 in 3D-HST (Skelton et al., 2014) in many surveys of the HUDF/CDFS region. It is listed as ID #884 in the catalog of the MUSE-HUDF mosaic by Inami et al. (2017), and for brevity, we will refer to it by only this number in the following. It is a luminous blue galaxy with $m_{\mathrm{AB}}=21.7$ in the ACS/F775W band, implying $L\approx L^{\star}$ at $z=0.737$. Figure 1 shows a deep HST image of the galaxy and its immediate surroundings within ±$\pm 11\arcsec$ ($\pm$80 kpc). Several of the galaxies close to the line of sight have redshifts measured with MUSE and are included in the Inami et al. (2017) sample; these are labeled in the figure. We also provide a few photometric redshifts (taken from Rafelski et al., 2015) where this may be of interest. Galaxies unlabelled in Fig. 1 have no reliable redshift information or are most likely unrelated to #884. The physically closest galaxy – marked as ‘NB1’ – is MUSE-HUDF #6516 ($m_{\mathrm{AB}}=23.8$) at 53 kpc projected distance and with only $-85$ km s-1 redshift difference. Other than this, there are no surrounding objects that are plausibly located nearby in real space; the galaxy at $z=0.734$ near the bottom of the field has a redshift separation of $-1000$ km s-1 and is thus probably already in the foreground. On transverse scales of up to $\sim 1$ Mpc, the full MUSE-HUDF Mosaic reveals 23 (11) galaxies within $\pm$ 1000 (300) km s-1, most of them faint emission line objects scattered all over the field. There is no evidence to support #884 being located in a cluster or rich group environment.
@@
Other projected neighbors to #884 are marked in Fig. 1 as ‘BG’ (background) and ‘FG’ (foreground). Since these galaxies are faint, we used the TDOSE software package (Three-Dimensional Optimal Spectral Extraction, Schmidt et al., 2019) to perform an optimal spectra extraction. Each of the galaxies was modelled in the HST data as a superposition of multiple components, which after blurring to the MUSE point spread function (PSF) were then used to define the spatial profiles of the galaxies used in the extraction. We briefly comment on three of these projected neighbors:
@@
  • BG1 (= MUSE-HUDF #6943 in Inami et al. 2017) is a spectroscopically confirmed system at $z=1.221$, at 4″ distance to #884, and with $m_{\mathrm{AB,F775W}}=24.5$ it is sufficiently bright to produce a well-detected continuum in the MUSE spectrum. The continuum-subtracted data cube shows two significantly negative peaks – indicative of absorption in the subtracted continuum – at the location of BG1 and at the wavelengths of the Mg II emission around #884. We analyze and discuss this interesting sight line in Sect. 8.3.
@@
  • The galaxy marked as FG1 ($m_{\mathrm{AB,F775W}}=25.3$) has a projected distance of only 2″ to #884, but no redshift in Inami et al. (2017). The photometric redshift estimate in Rafelski et al. (2015) puts it in the foreground, $0.530.38<z_{\mathrm{BPZ}}<$ 0.53. The extracted MUSE spectrum does not show any emission features, apart from [O II] spilling over from #884, and it is too noisy to allow for a confident identification of individual absorption lines. It, however, agrees well in shape with a synthetic galaxy spectrum from the Jimenez et al. (2004) spectral library for an age of $\sim 2$ Gyr at a redshift of $z=0.621\pm 0.002$. This redshift is outside the range estimated for the photometric redshift. However, assessing the reliability of the Rafelski et al. (2015) catalog is outside the scope of this paper. While our redshift measurement is not as secure as those based on multiple individual spectral lines, we consider it sufficiently robust to conclude that this galaxy is indeed located in the foreground of #884 and thus cannot provide another absorption line of sight.